OpenStudy (anonymous):

what is Lim cos(((3pi/2)+h)-cos(3pi/2))/h as h approaches 0

OpenStudy (kc_kennylau):

\[\Large\lim_{h\rightarrow0}\frac{\cos\left(\frac{3\pi}2+h\right)-\cos\left(\frac{3\pi}2\right)}h\]

OpenStudy (kc_kennylau):

\[\Huge=\lim_{h\rightarrow0}\Large\frac{\cos\frac{3\pi}2\cos h-\sin\frac{3\pi}2\sin h-\cos\frac{3\pi}2}h\]

OpenStudy (kc_kennylau):

\[\Huge=\lim_{h\rightarrow0}\Large\frac{\sin h}h\]

OpenStudy (kc_kennylau):

=1

OpenStudy (anonymous):

Thank you!

OpenStudy (kc_kennylau):

no problem :)